# Arboreal Galois groups of postcritically finite quadratic polynomials

### v1

@@ 1.1. Arboreal Galois groups
contains no critical values of $f$, then each $f^{-n}(x_{0})$ has exactly $2^{n}$ elements. If, in addition, $x_{0}$ is not periodic under $f$, then the sets $f^{-n}(x_{0})$ are pairwise disjoint, and hence $\operatorname{Orb}_{f}^{-}(x_{0})$ has the structure of an infinite binary rooted tree $T_{\infty}$, with $x\in f^{-(n+1)}(x_{0})$ connected to $f(x)\in f^{-n}(x_{0})$ by an edge. Thus, the action of the Galois group $G_{\infty}$ on the backward orbit induces an embedding of $G_{\infty}$ into the automorphism group $\operatorname{Aut}(T_{\infty})$ of the tree. Similarly, for each $n\geq 0$, the action of $G_{n}$ on $f^{-n}(x_{0})$ induces an embedding of $G_{n}$ into the automorphism group $\operatorname{Aut}(T_{n})$ of the finite binary rooted tree $T_{n}$ with $n$ levels. For this reason, the groups $G_{n}$ and $G_{\infty}$ have come to be known as arboreal Galois groups. Moreover, given our interest in this action, whenever we discuss homomorphisms or isomorphisms between groups acting on trees, we always mean homomorphisms that are equivariant with respect to this action. We note that the problem of fully understanding the arboreal Galois groups has generated a great deal of research in the recent years; see [ABC+22, BDG+21, BFH+17, BGJT23, BJ19, JKMT16, Juu19, Pin13], for example.
@@ 1.4. Our approach
This paper is the first of a series of two papers in which we have two main goals. First, for each pair of integers $r>s\geq 0$, we construct subgroups $B_{r,s,\infty}\subseteq M_{r,s,\infty}$ of $\operatorname{Aut}(T_{\infty})$, coinciding with Pink’s group $G^{\textup{Pink}}_{r,s,\infty}\cong G^{\textup{geom}}\subseteq G^{\textup{arith}}$, and we show that the arboreal Galois group $G_{\infty}$ is isomorphic to a subgroup of $M_{r,s,\infty}$. Our arguments apply over general fields with arbitrary base points, rather than restricting to the case $K=k(t)$ with base point $t$. Our approach to this problem is also more concrete than that of Pink; the groups $B_{r,s,\infty}$ and $M_{r,s,\infty}$ are defined not by generators but rather as the set of all $\sigma\in\operatorname{Aut}(T_{\infty})$ satisfying certain parity conditions, which are also used to describe how elements of $G_{\infty}$ act on the roots of unity contained in $K_{\infty}$. One advantage of this approach is that it can allow us to describe the intersections $K_{n}\cap k(\mu_{2^{\infty}})$ with a great deal of precision (see Corollary 6.5).
@@
This paper handles the periodic case $s=0$ for arbitrary $r\geq 1$, for which we denote the above groups simply as $M_{r,\infty}$ and $B_{r,\infty}$ (see Definition 1.3). We handle the strictly preperiodic cases $s>0$ in a separate paper; nevertheless, even though additional technical complications arise in the strictly preperiodic cases, the main ideas for all of our constructions already arise in the periodic case considered here.
@@ 1.6. Statement of our main result
When $k$ contains $i$, the conditions of Theorem 1.4 can never hold, as condition (1) necessarily fails. However, our methods can still be used to prove a slightly more complicated result involving an appropriate subgroup of $M_{r,\infty}$. Specifically, this subgroup is the inverse image under $P_{r}$ of the image in $\mathbb{Z}_{2}^{\times}$ of the $2$-adic cyclotomic character of $\operatorname{Gal}({\bar{k}}/k)$.
@@ 1.7. Outline of the paper
Section 2 concerns a useful elementary result for general quadratic polynomials that underlies many of our subsequent arguments. In Section 3 we present explicit formulas yielding $2$-power roots of unity as arithmetic combinations of preimages of an arbitrary root point under our quadratic polynomial $f(z)=z^{2}+c$ when the critical point is periodic. Section 4 is devoted to proving that $M_{r,\infty}$ is a subgroup of $\operatorname{Aut}(T_{\infty})$ and that $B_{r,\infty}$ is simply the kernel of $P_{r}$ (this is done in Theorem 4.1), while Section 5 shows how $M_{r,\infty}$ and $B_{r,\infty}$ realize the arboreal Galois action. In Section 6, we prove that our group $B_{r,\infty}$ concides with Pink’s group $G^{\textup{geom}}$, and that Pink’s larger group $G^{\textup{arith}}$ is contained in our $M_{r,\infty}$. Finally, in Section 7, we prove Theorem 1.4, giving necessary and sufficient conditions for the Galois group $G_{\infty}$ to be the full group $M_{r,\infty}=G^{\textup{arith}}$.
@@ 3. Roots of unity arising in backward orbits
In the statement of our next result, recall Definition 1.1 of the set of words $W(r,i)$ of length $ri-1$ satisfying a certain restriction.
@@ 4. A preliminary result regarding the associated arboreal subgroup
Step 3. For $\sigma,\tau,x$ as in Step 2, a straightforward induction on $i\geq 0$ gives
@@
$P_{r}(\sigma)P_{r}(\tau,x)$ $=(-1)^{\operatorname{Par}(\sigma\tau,x)}+Q(\sigma\tau,xb)-Q(\sigma\tau,xa)+Z_{r}(\sigma,\tau,xb)-Z_{r}(\sigma,\tau,xa)$
@@
Step 4. We now show that $M_{r,\infty}$ is a subgroup of $\operatorname{Aut}(T_{\infty})$. The identity $e\in\operatorname{Aut}(T_{\infty})$ clearly satisfies $Q_{r}(e,x)=0$ and $\operatorname{Par}(e,x)=0$ for all nodes $x$, whence $P(e,x)=1$, so that $e\in M_{r,\infty}$. Given $\sigma,\tau\in M_{r,\infty}$ and $x_{1},x_{2}\in X$, we have
@@ 5. The action of Galois on roots of unity
The following result shows that the group $M_{r,\infty}$ is determined solely by the restriction that a Galois element $\sigma$ must act consistently on every instance of any root of unity $\zeta_{2^{n}}$.
@@
We first claim for all $i\geq 0$ and any finite word $w$ (in the alphabet $\{a,b\}$), we have
@@
where for $i=0$, we interpret this equation as saying $\sigma(wa)=\zeta_{2}^{-\operatorname{Par}(\sigma,w)}\sigma(w)a$. When $\operatorname{Par}(\sigma,w)=1$, we have $\sigma(wa)=\sigma(w)b$, so equation (14) follows from Lemma 3.2 applied to $\sigma(w)$. On the other hand, when $\operatorname{Par}(\sigma,w)=0$, we have $\sigma(wa)=\sigma(w)a$ and hence the set of words in product on the left of the equation is the same as that on the right hand side, so equation (14) is vacuously true in this case.
@@
Write $S_{i}:={\sum_{w\in W(r,i)}\operatorname{Par}(\sigma,xaw)}$. Note, any $w\in W_{(r,j)}$ can be written as $w_{1}aw_{2}a\dots w_{j}$ for $w_{1},\dots,w_{j}\in\{a,b\}^{r-1}$. Then alternately applying equation (14) and equation (15), we have
@@ 6.1. The geometric Galois group
for all infinite words $w\in\{a,b\}^{\mathbb{N}}$. In addition, for any $\sigma_{a},\sigma_{b}\in\operatorname{Aut}(T_{\infty})$, the automorphism $(\sigma_{a},\sigma_{b})$ is the element $\sigma\in\operatorname{Aut}(T_{\infty})$ given by
@@ 6.2. The arithmetic Galois group
By Theorem 5.1, there is an equivariant injective homomorphism $\rho:G^{\textup{arith}}\hookrightarrow M_{r,\infty}$. Moreover, by Theorem 6.2, restricting this homomorphism to $\operatorname{Gal}(K_{\infty}/k_{\infty}(t))\cong G^{\textup{geom}}$ yields an isomorphism $G^{\textup{geom}}\cong B_{r,\infty}$.
@@
If $[k(\zeta_{8}):k]=4$, then $K_{n}\cap\overline{k}=k(\zeta_{2^{e}})$ for all $n\geq 1$, where $e=\lfloor\frac{n-1}{r}\rfloor+1$. In particular, $\zeta_{2^{n}}\in K_{rn+1}$, and if $n\geq 2$, then $\zeta_{2^{n}}\notin K_{rn}$.
@@ 7. Obtaining the arboreal Galois groups
Let $k$ be any field with $[k(\zeta_{8}):k]=4$. Suppose $f(z)=z^{2}+c\in k[z]$ is PCF. Let $K=k(t)$, let $x_{0}=t$, and let $K_{\infty}$ be the resulting arboreal extension of $K$. Let $A\in k$ such that $\sqrt{A}\in K_{\infty}$. Then $\sqrt{A}\in k(\zeta_{8})$.
@@
Since $\mathbb{Z}_{2}^{\times}$ has only three index $2$ subgroups, $k_{\infty}$ contains only three quadratic extensions of $k$ (formed by adjoining one of $\sqrt{-1}$, $\sqrt{2}$, or $\sqrt{-2}$), all of which are contained in $k(\zeta_{8})$, there are similarly only three quadratic extensions of $k$ in $k(\mu_{2^{\infty}})$, all contained in $k(\zeta_{8})$. Thus, $\sqrt{A}\in k(\zeta_{8})$. ∎
The following argument has been presented in various degrees of generality in [Odo88, Lemma 4.2], [Sto92, Lemmas 1.5 and 1.6], [Jon13, Section 2.2], [BD23, Proposition 5.3, Theorem 6.5]. We include a proof here for completeness.
@@
Then the following are equivalent.
@@
Thus, we have $\Delta_{i}=c_{i}^{2}D_{i}$ for some $c_{i}\in K$, so $\Delta_{i}$ is a square in $K$. Therefore, every $\sigma\in G_{i}$ acts as an even permutation of the elements of $f^{-i}(x_{0})$, so that $G_{i}\not\cong\operatorname{Aut}(T_{i})$, and hence $G_{n}\not\cong\operatorname{Aut}(T_{n})$ for all $n\geq i$.
@@
to $K_{i-1}$. By Kummer theory, the degree $[K_{i}:K_{i-1}]$ is the order of the group generated by the classes of the elements $\beta_{j}-c$ in $K_{i-1}^{\times}/(K_{i-1}^{\times})^{2}$. This order is $2^{2^{i-1}}/\#V$, where
$V=\Big{\{}(\epsilon_{1},\dots,\epsilon_{2^{i-1}})\in\mathbb{F}_{2}^{2^{i-1}}:\prod_{j}(b_{j}-c)^{\epsilon_{j}}\in(K_{i-1}^{\times})^{2}\Big{\}}.$
@@
The implications (5$\Leftrightarrow$(4$\Rightarrow$(3$\Leftrightarrow$(2) are trivial. Thus, it suffices to show (1$\Rightarrow$(5) and (2$\Rightarrow$(1). Define
@@
First assume statement (1), that $[L(\zeta_{8}):K]=2^{r+2}$. Then we have $[k(\zeta_{8}):k]=4$ and $[L(\zeta_{8}):k(\zeta_{8})]=2^{r}$. As at the start of Section 6, let $k_{\infty}:=K_{t,\infty}\cap\overline{k}$ be the constant field extension in the setting of the function field $k(t)$ with root point $t$ in place of $x_{0}$. Let $k^{\prime}$ denote the compositum of the degree $2$ extensions of $k$ contained in $k_{\infty}$. By Lemma 7.1, we have $k^{\prime}=k(\zeta_{8})$. Moreover, since $[L(\zeta_{8}):k^{\prime}]=2^{r}$, Lemma 7.2 implies that
@@
Noting that the forward orbit of 00 has cardinality $r$, we now apply [BGJT23, Theorem 4.6] (whose hypotheses assumes $k$ is a number field, but only to ensure that $[k^{\prime}:k_{1}]$ is finite, a fact which is evident in our case). This result says, given the length $r$ of the forward orbit of 00, along with condition (23), that $G_{x_{0},\infty}\cong M_{r,\infty}$, i.e., that statement (5) holds.
Finally, assume statement (2), which implies both that $|G_{x_{0},r}|=|M_{r,r}|=|\operatorname{Aut}(T_{r})|$ and that $[K_{x_{0},r}(\zeta_{8}):K_{x_{0},r}]=4$. Since $|G_{x_{0},r}|=|\operatorname{Aut}(T_{r})|$, Lemma 7.2 implies that $[L:k]\geq 2^{r}$. On the other hand, because $L\subseteq K_{x_{0},r}$, we have $[L(\zeta_{8}):L]\geq[K_{x_{0},r}(\zeta_{8}):K_{x_{0},r}]=4$. Therefore,
$2^{r+2}\geq[L(\zeta_{8}):k]=[L(\zeta_{8}):L]\cdot[L:k]\geq 2^{r+2},$

### v2

@@ 1.1. Arboreal Galois groups
contains no critical values of $f$, then each $f^{-n}(x_{0})$ has exactly $2^{n}$ elements. If, in addition, $x_{0}$ is not periodic under $f$, then the sets $f^{-n}(x_{0})$ are pairwise disjoint, and hence $\operatorname{Orb}_{f}^{-}(x_{0})$ has the structure of an infinite binary rooted tree $T_{\infty}$, with $x\in f^{-(n+1)}(x_{0})$ connected to $f(x)\in f^{-n}(x_{0})$ by an edge. Thus, the action of the Galois group $G_{\infty}$ on the backward orbit induces an embedding of $G_{\infty}$ into the automorphism group $\operatorname{Aut}(T_{\infty})$ of the tree. Similarly, for each $n\geq 0$, the action of $G_{n}$ on $f^{-n}(x_{0})$ induces an embedding of $G_{n}$ into the automorphism group $\operatorname{Aut}(T_{n})$ of the finite binary rooted tree $T_{n}$ with $n$ levels. For this reason, the groups $G_{n}$ and $G_{\infty}$ have come to be known as arboreal Galois groups. Moreover, given our interest in this action, whenever we discuss homomorphisms or isomorphisms between groups acting on trees, we always mean homomorphisms that are equivariant with respect to this action. We note that the problem of fully understanding the arboreal Galois groups has generated a great deal of research in the recent years; see [ABC+22, BDG+21, BFH+17, BGJT25, BHL17, BJ19, JKMT16, Juu19, Pin13], for example.
@@ 1.4. Our approach
This paper is the first of a pair of papers in which we have two main goals. First, for each pair of integers $r>s\geq 0$, we construct subgroups $B_{r,s,\infty}\subseteq M_{r,s,\infty}$ of $\operatorname{Aut}(T_{\infty})$, coinciding with Pink’s groups $G^{\textup{Pink}}_{r,s,\infty}\cong G^{\textup{geom}}\subseteq G^{\textup{arith}}$, and we show that the arboreal Galois group $G_{\infty}$ is isomorphic to a subgroup of $M_{r,s,\infty}$. Our arguments apply over general fields with arbitrary base points, rather than restricting to the case $K=k(t)$ with base point $t$. Our approach to this problem is also more concrete than that of Pink; the groups $B_{r,s,\infty}$ and $M_{r,s,\infty}$ are defined not by generators but rather as the set of all $\sigma\in\operatorname{Aut}(T_{\infty})$ satisfying certain parity conditions, which are also used to describe how elements of $G_{\infty}$ act on the roots of unity contained in $K_{\infty}$. One advantage of this approach is that it can allow us to describe the intersections $K_{n}\cap k(\mu_{2^{\infty}})$ with a great deal of precision (see Corollary 6.5).
@@
This paper handles the periodic case $s=0$ for arbitrary $r\geq 1$, for which we denote the above groups simply as $M_{r,\infty}$ and $B_{r,\infty}$ (see Definition 1.4). We handle the strictly preperiodic cases $s>0$ in a separate paper [BGJT]; nevertheless, even though additional technical complications arise in the strictly preperiodic cases, the main ideas for all of our constructions already arise in the periodic case considered here.
Remark 1.1.
More generally, for any $d\geq 2$, the unicritical polynomial $f(z)=z^{d}+c$ is PCF if and only if the critical point at 00 is preperiodic. There has been growing interest in the associated arithmetic dynamics, for example in [BDG+21, Buf18, BEK22, Gok20, HT15], including arboreal Galois investigations in [BGJT25, BHL17]. In particular, in [BGJT25], we considered the constant field extensions $K_{n}\cap\overline{k}$ for PCF maps $f(z)=z^{p}+c$ with $p$ prime, as well as the relationship between the generic arboreal Galois groups (when $K=k(t)$ and $x_{0}=t$) and the corresponding groups when specializing the parameter $t$. It would be interesting to find explicit descriptions of the of the resulting arboreal Galois groups for various PCF orbit structures for $z^{d}+c$. For now, however, we restrict our attention to the case $d=2$, affording us both Pink’s work and the relative simplicity of binary trees versus $d$-ary trees.
@@ 1.6. Statement of our main result
When $K$ contains $\sqrt{-1}$, the conditions of Theorem 1.5 can never hold, as condition (1) necessarily fails. However, our methods can still be used to prove a slightly more complicated result involving an appropriate subgroup of $M_{r,\infty}$. Specifically, this subgroup is the inverse image under $P_{r}$ of the image in $\mathbb{Z}_{2}^{\times}$ of the $2$-adic cyclotomic character of $\operatorname{Gal}({\bar{K}}/K)$.
@@ 1.7. Outline of the paper
Section 2 concerns a useful elementary result for general quadratic polynomials that underlies many of our subsequent arguments. In Section 3 we present explicit formulas yielding $2$-power roots of unity as arithmetic combinations of preimages of an arbitrary root point under our quadratic polynomial $f(z)=z^{2}+c$ when the critical point is periodic. Section 4 is devoted to proving that $M_{r,\infty}$ is a subgroup of $\operatorname{Aut}(T_{\infty})$ and that $P_{r}$ is a group homomorphism with kernel $B_{r,\infty}$ (this is done in Theorem 4.1), while Section 5 shows how $M_{r,\infty}$ and $B_{r,\infty}$ realize the arboreal Galois action. In Section 6, we prove that our group $B_{r,\infty}$ concides with Pink’s group $G^{\textup{geom}}$, and that Pink’s larger group $G^{\textup{arith}}$ is contained in our $M_{r,\infty}$. (Recall that $G^{\textup{geom}}$ and $G^{\textup{arith}}$ are special cases of $G_{\infty}$ for $K=\overline{k}(t)$ and $K=k(t)$, respectively, with $x_{0}=t$.) Finally, in Section 7, we prove Theorem 1.5, giving necessary and sufficient conditions for the Galois group $G_{\infty}$ to be the full group $M_{r,\infty}$.
@@ 3. Roots of unity arising in backward orbits
In the statement of our next result, recall Definition 1.2 of the set $W(r,i)$ of words $s_{1}\cdots s_{ri-1}$ of length $ri-1$ with $s_{j}=a$ whenever $r|j$. In addition, observe that the nodes of $T_{\infty}$ with labels $[xawa]$ for $w\in W(r,i)$ form a choice of a set $A_{i+1}$ as in Lemma 3.1, and those with labels $[xbwa]$ form a choice of a set $B_{i+1}$.
@@ 4. A preliminary result regarding the associated arboreal subgroup
Step 3. We will now show that $Z_{r}$ is indeed identically zero, and furthermore, with $\sigma$ and $\tau$ as in Step 2, that $P_{r}(\sigma)P_{r}(\tau,x)=P_{r}(\sigma\tau,x)$.
For $\sigma,\tau,x$ as in Step 2, a straightforward induction on $i\geq 0$ gives
@@
$P_{r}(\sigma)P_{r}(\tau,x)$ $=(-1)^{\operatorname{Par}(\sigma\tau,x)}+Q_{r}(\sigma\tau,xb)-Q_{r}(\sigma\tau,xa)+Z_{r}(\sigma,\tau,xb)-Z_{r}(\sigma,\tau,xa)$
@@
Step 4. We now show that $M_{r,\infty}$ is a subgroup of $\operatorname{Aut}(T_{\infty})$. The identity $e\in\operatorname{Aut}(T_{\infty})$ clearly satisfies $Q_{r}(e,x)=0$ and $\operatorname{Par}(e,x)=0$ for all nodes $x$, whence $P_{r}(e,x)=1$, so that $e\in M_{r,\infty}$. Given $\sigma,\tau\in M_{r,\infty}$ and $x_{1},x_{2}\in X$, we have
@@ 5. The action of Galois on roots of unity
Our next result shows that $P_{r}(\sigma,x)$ determines the action of $\sigma\in\operatorname{Aut}(T_{\infty})$ on the roots of unity $\zeta_{2^{n}}$ constructed in Lemma 3.2, thus explaining the otherwise mysterious Definition 1.3 of $P_{r}(\sigma,x)$. In addition, a Galois element $\sigma$ must of course act consistently on roots of unity, thus explaining the defining condition (5) of $M_{r,\infty}$, that $P_{r}(\sigma,x)$ be independent of $x$. That is, for $\sigma\in\operatorname{Aut}(T_{\infty})$ to belong to $M_{r,\infty}$, Definition 1.4 imposes only the rudimentary condition that $\sigma$ must act consistently on roots of unity $\zeta_{2^{n}}$.
@@
We first claim for all $i\geq 0$ and any word $w$ (in the alphabet $\{a,b\}$), we have
@@
where for $i=0$, we interpret this equation as saying $[\sigma(wa)]=\zeta_{2}^{-\operatorname{Par}(\sigma,w)}[\sigma(w)a]$. When $\operatorname{Par}(\sigma,w)=1$, we have $\sigma(wa)=\sigma(w)b$, so equation (14) follows from Lemma 3.2 applied to $\sigma(w)$. On the other hand, when $\operatorname{Par}(\sigma,w)=0$, we have $\sigma(wa)=\sigma(w)a$ and hence the set of words in product on the left of the equation is the same as that on the right hand side, so equation (14) is vacuously true in this case.
@@
Write $S_{i}:={\sum_{w\in W(r,i)}\operatorname{Par}(\sigma,xaw)}$. Note, any $w\in W(r,j)$ can be written as $w_{1}aw_{2}a\dots w_{j}$ for $w_{1},\dots,w_{j}\in\{a,b\}^{r-1}$. Then alternately applying equation (14) and equation (15), we have
@@ 6.1. The geometric Galois group
for all words $w\in\{a,b\}^{\mathbb{N}}$. In addition, for any $\sigma_{a},\sigma_{b}\in\operatorname{Aut}(T_{\infty})$, the automorphism $(\sigma_{a},\sigma_{b})$ is the element $\sigma\in\operatorname{Aut}(T_{\infty})$ given by
@@ 6.2. The arithmetic Galois group
By Theorem 5.1, there is an injective homomorphism $\rho:G^{\textup{arith}}\hookrightarrow M_{r,\infty}$. Moreover, by Theorem 6.2, restricting this homomorphism to $\operatorname{Gal}(K_{\infty}/k_{\infty}(t))\cong G^{\textup{geom}}$ yields an isomorphism $G^{\textup{geom}}\cong B_{r,\infty}$.
@@
If $[k(\zeta_{8}):k]=4$, then $K_{n}\cap\overline{k}=k(\zeta_{2^{e}})$ for all $n\geq 1$, where $e=\lfloor\frac{n-1}{r}\rfloor+1$. In particular, $\zeta_{2^{n}}\in K_{r(n-1)+1}$, and if $n\geq 2$, then $\zeta_{2^{n}}\notin K_{r(n-1)}$.
@@ 7. Obtaining the arboreal Galois groups
Let $k$ be any field with $[k(\zeta_{8}):k]=4$. Suppose $f(z)=z^{2}+c\in k[z]$ with 00 periodic. Let $K=k(t)$, let $x_{0}=t$, and let $K_{\infty}$ be the resulting arboreal extension of $K$. Let $A\in k$ such that $\sqrt{A}\in K_{\infty}$. Then $\sqrt{A}\in k(\zeta_{8})$.
@@
Since $\mathbb{Z}_{2}^{\times}$ has only three index $2$ subgroups, $k_{\infty}$ contains only three quadratic extensions of $k$ (formed by adjoining one of $\sqrt{-1}$, $\sqrt{2}$, or $\sqrt{-2}$), all of which are contained in $k(\zeta_{8})$; therefore, $\sqrt{A}\in k(\zeta_{8})$. ∎
The following argument has been presented in various degrees of generality in [Odo88, Lemma 4.3], [Sto92, Lemmas 1.5 and 1.6], [Jon13, Section 2.2], [BD24, Proposition 5.3, Theorem 6.5]. We include a proof here for completeness.
@@
Then for $n\geq 1$, the following are equivalent:
@@
Thus, we have $\Delta_{i}=c_{i}^{2}D_{i}$ for some $c_{i}\in K$, so $\Delta_{i}$ is a square in $K\left(\sqrt{D_{1}},\dots,\sqrt{D_{i-1}}\right)$. Therefore, every $\sigma\in G_{i}$ acts as an even permutation in ${\rm Gal}(K_{i}/K_{i-1})$; hence, $G_{i}\not\cong\operatorname{Aut}(T_{i})$, and so, $G_{n}\not\cong\operatorname{Aut}(T_{n})$ for all $n\geq i$.
@@
to $K_{i-1}$. By Kummer theory, the degree $[K_{i}:K_{i-1}]$ is the order of the group generated by the classes of the elements $\beta_{j}-c$ in $K_{i-1}^{\times}/(K_{i-1}^{\times})^{2}$. This order is $2^{2^{i-1}}/|V|$, where
$V=\Big{\{}(\epsilon_{1},\dots,\epsilon_{2^{i-1}})\in\mathbb{F}_{2}^{2^{i-1}}:\prod_{j}(\beta_{j}-c)^{\epsilon_{j}}\in(K_{i-1}^{\times})^{2}\Big{\}}.$
@@
The implications (5$\Leftrightarrow$(4$\Rightarrow$(3$\Leftrightarrow$(2) are trivial. Thus, it suffices to show (1$\Rightarrow$(5) and (3$\Rightarrow$(1). Define
@@
First assume statement (1), that $[L(\zeta_{8}):K]=2^{r+2}$. Then we have $[K(\zeta_{8}):K]=4$ and $[L(\zeta_{8}):K(\zeta_{8})]=2^{r}$. Consider the arboreal extension $K(t)_{t,\infty}$ of the function field $K(t)$ for the root point $t$ in place of $x_{0}$, and define $K^{\prime}$ to be the compositum of the degree $2$ extensions of $K$ contained in $K(t)_{t,\infty}$. By Lemma 3.1, we have $\zeta_{8}\in K(t)_{t,\infty}$, and hence $\zeta_{8}\in K^{\prime}$. Therefore, by Lemma 7.1, we in fact have $K^{\prime}=K(\zeta_{8})$.
Since we now have $[L(\zeta_{8}):K^{\prime}]=2^{r}$, Lemma 7.2 implies that
@@
Noting that the forward orbit of 00 has cardinality $r$, we now apply [BGJT25, Theorem 4.6] (whose hypotheses assume $K$ is a number field, but only to ensure that $[K^{\prime}:K]$ is finite, a fact which is evident in our case). This result says, given the length $r$ of the forward orbit of 00, along with condition (23), that $G_{x_{0},\infty}\cong M_{r,\infty}$, i.e., that statement (5) holds.
Finally, assume statement (3), which implies that $G_{x_{0},r}\cong M_{r,r}\cong\operatorname{Aut}(T_{r})$, and hence, by Lemma 7.2, that $[L:K]\geq 2^{r}$.
We claim that statement (3) also implies $[K_{x_{0},r}(\zeta_{8}):K_{x_{0},r}]=4$. To see this, define $H\subseteq M_{r,2r+1}$ to be the subgroup consisting of all $\sigma\in M_{r,2r+1}$ fixing level $r$ of the tree, and let $\tilde{H}\subseteq G_{x_{0},2r+1}$ be the image of $H$ under the assumed isomorphism $M_{r,2r+1}\cong G_{x_{0},2r+1}$. Combining Theorem 6.4 for $k=\mathbb{Q}(c)$ with Corollary 6.5, the orbit of $\zeta_{8}$ under $H\subseteq M_{r,2r+1}$ must have cardinality 4. (Recall that this action of $M_{r,\infty}$ on $\zeta_{8}$ is given by Theorem 5.1.) Since $\zeta_{8}\in K_{x_{0},2r+1}$ by Lemma 3.1, the orbit of $\zeta_{8}$ under $\tilde{H}\subseteq G_{x_{0},2r+1}$ must also have cardinality 4. But $\tilde{H}=\operatorname{Gal}(K_{x_{0},2r+1}/K_{x_{0},r})$, and hence $[K_{x_{0},r}(\zeta_{8}):K_{x_{0},r}]=4$, proving our claim.
Because $L\subseteq K_{x_{0},r}$, it follows that $[L(\zeta_{8}):L]\geq 4$. Thus, recalling our observation above that $[L:K]\geq 2^{r}$, we have
$2^{r+2}\geq[L(\zeta_{8}):K]=[L(\zeta_{8}):L]\cdot[L:K]\geq 4\cdot 2^{r}=2^{r+2},$
@@
Acknowledgments. The first author gratefully acknowledges the support of NSF grant DMS-2401172. We thank the anonymous referees for their useful comments and suggestions.